OpenStudy (anonymous):

parallel lines "L" and "M" are cut by a transversal. Angle "A" measures ( 15x + 6 ) degrees and angle "F" measures (5x + 14) degrees. What are the degree measures of angles "C" and "H" ?

OpenStudy (heisenberg):

could you upload an image? i don't know what C and H are.

OpenStudy (anonymous):

they are the angles. here. ignore the bottom part

OpenStudy (caboose):

C+H=180 degrees, solve for x, plug into equations, get degrees for each angle

OpenStudy (anonymous):

this is what i did... 15x + 6 = 5x+ 14

OpenStudy (anonymous):

is that how i am suppose to set it up?

OpenStudy (caboose):

The equation you are solving should be (15x+6)+(5x+14)=180

OpenStudy (anonymous):

i missed out on that.

OpenStudy (anonymous):

combine like terms right?

OpenStudy (caboose):

yes

OpenStudy (anonymous):

20x + 20 = 180

OpenStudy (caboose):

mhmm, then move the 20 to the other side.

OpenStudy (anonymous):

x = 8

OpenStudy (anonymous):

then i wud plug it in

OpenStudy (caboose):

correct

OpenStudy (anonymous):

dude. thank you so much man!

OpenStudy (anonymous):

i got my answers. thank you sooo soo much! you made it clearer to me!

OpenStudy (caboose):

No problem, glad i could help. I've been there haha

OpenStudy (anonymous):

lol. thank you! =)